### v1

# Log-concavity And The Multiplicative Properties of Restricted Partition Functions

Abstract.
The partition function $p(n)$ and many of its related restricted partition functions have recently been shown independently to satisfy log-concavity: $p(n)^{2}\geq p(n-1)p(n+1)$, and satisfy the inequality: $p(n)p(m)\geq p(n+m)$ with only finitely many instances of equality or failure. This paper proves that this is no coincidence, that any log-concave sequence $\{x_{n}\}$ satisfying a particular initial condition likewise satisfies the inequality $x_{n}x_{m}\geq x_{n+m}$. This paper further determines that these conditions are sufficient but not necessary and considers various examples to illuminate the situation.
Key words and phrases:
Partition function, log-concavity, log-convexity, Higher Order Turàn inequalities, restricted partition functions
2020 Mathematics Subject Classification:
05A17; 11P82; 11B99
1. Introduction
The partition function $p(n)$ enumerates the number of partitions of a positive integer $n$ where the partitions are positive integer sequences $\lambda=(\lambda_{1},\lambda_{2},...)$ with $0<\lambda_{1}\leq\lambda_{2}\leq~{}\dots$ and $\sum_{j\geq 1}{\lambda_{j}}=n$. For example, $p(4)=5$ since the only ways to partition $4$ are $4$, $3+1$, $2+2$, $2+1+1$, and $1+1+1+1$. Recent interest regarding the various properties of $p(n)$ led to a conjecture by Chen [7], and subsequent proof by DeSalvo & Pak [10], that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ is log-concave for $n>25$. A sequence of integers $\{x_{n}\}_{n\in\mathbb{N}}$ is log-concave if $x_{n}^{2}~{}\geq~{}x_{n-1}x_{n+1}$ and is is log-convex if $x_{n}^{2}\leq x_{n-1}x_{n+1}$ for all $n$.
Log-concavity had been previously established in 1978 by Nicholas [29] for the partition function, but there was a revitalization of interest following the result of DeSalvo & Pak [10]. New results found in the work of Bessenrodt & Ono [30] in 2014 proved that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $p(n)p(m)\geq p(n+m)$ where instances of equality occur at $(2,6),(3,4)$, and instances of failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(5,5)$ and nowhere else.
For convenience, this paper uses the notation: a sequence $\{x_{n}\}$ is multiplicatively abundant (or simply abundant) if $x_{n}x_{m}\geq x_{n+m}$, and is multiplicatively deficient (or simply deficient) if $x_{n}x_{m}\leq x_{n+m}$ for all but finitely many $n$ and $m$. In these terms, Bessenrodt & Ono [30] proved that $\{p(n)\}$ is abundant. Proofs that the partition function is log-concave or abundant have so far relied on an analytic result of Rademacher type due to Lehmer [26] until 2017 when A. Alanazi, S.M. Gagola III, & A.O. Munagi [2] published a combinatorial proof that $\{p(n)\}$ is abundant.
Many sequences (especially sequences arising from restricted partitions) are sporadic for small indices but become and stay log-concave for sufficiently large indices. The main goal of this paper is to prove that any sequence that is eventually log-concave is also eventually abundant, provided the sequence also satisfies a necessary (but not sufficient) initial condition.
Theorem 1.1.
Let $\{x_{n}\}$ be a sequence of positive real numbers and let $N$ be a positive integer such that for all $n>N$, the sequence satisfies $x_{n}^{2}\geq x_{n-1}x_{n+1}$. Then for all $n>N$ and $m\geq 0$, the sequence $\{x_{n}\}$ satisfies:
$(x_{N})^{\frac{m}{n-N+m}}(x_{n+m})^{\frac{n-N}{n-N+m}}\leq x_{n}\leq x_{N}\left(\frac{x_{N+1}}{x_{N}}\right)^{n-N}.$ (1)
Moreover, if
$(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$ (2)
then $x_{n+m}\leq x_{n}x_{m}$ for all $n,m>N$.
Note that if $\{x_{n}\}$ is log-concave for all $n\geq 1$, then
$x_{1}^{\frac{m}{n+m-1}}(x_{n+m})^{\frac{n-1}{n+m-1}}\leq x_{n}\leq x_{1}\left(\frac{x_{2}}{x_{1}}\right)^{n-1}$
for all non-negative integer $m$. In light of Theorem 1.1 and because log-concavity has been established for $n>25$ [10], to show the partition function is abundant it suffices to show that Condition (2) holds for $p(26)=2436$ and $p(27)=3010$: it is easy to verify that in fact, $2436^{\frac{1}{26}}>3010^{\frac{1}{27}}$. An example demonstrates that Condition (2) in Theorem 1.1, for a log-concave sequence to be abundant, is required. If a log-concave sequence does not satisfy $(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\leq x_{n}x_{m}$ for all $n,m\geq N+1$.
Example 1. Define the sequence $\{a_{n}\}$ by $a_{n}=2^{n+1}-1$ for $n=0,1,2,\ldots$
$a_{n}=1,3,7,15,31,63,127,255,511,1023,2047,4095,8191,16383,32767,65535,\ldots$
The sequence is strictly log-concave:
$12^{2n+2}-2^{n+2}+1=(2^{n+1}-1)^{2}>(2^{n}-1)(2^{n+2}-1)=2^{2n+2}-2^{n}-2^{n+2}+$ 1
However, the sequence $\{(2^{n+1}-1)^{\frac{1}{n+1}}\}$ for $n=0,1,2,\ldots$, is a strictly increasing sequence whose limit equals $2$ as $n$ tends to infinity. Hence, $a(n)$ does not satisfy Condition (2):
$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}.$
In this particular case, $\{a_{n}\}$ is also not abundant (in fact, $\{a_{n}\}$ is deficient):
$(2^{n+1}-1)(2^{m+1}-1)=2^{n+m+2}-2^{n}-2^{m}+1<2^{n+m}-1$
for $n,m\geq 1$.
The interaction between log-behavior and multiplicative behavior in the setting of general sequences was been studied in the early 2000’s by Asai, Kubo, & Kuo [3] and by Došlić [12] who collectively established inequalities for Bell numbers, Motzkin numbers, Fine numbers, Fresnel numbers, Apèry numbers, large Schröder numbers, central Delannoy numbers, and Baxter permutations. Both Asai, Kubo, & Kuo [3] and Došlić [12] give a proof that any log-convex sequence is deficient, provided $x_{0}=1$. More recently, a similar version of this proof emerged in 2023 by Gajdzica, Miska, & Ulas [17] with the slight modification that $x_{0}\geq 1$. Unfortunately, there is some misunderstanding that every log-convex sequence is deficient. Note that the reciprocal of a log-concave sequence of positive numbers is a log-convex sequence, and vice versa.
$x_{n}^{2}\geq x_{n-1}x_{n+1}\qquad\iff\qquad\left(\frac{1}{x_{n}}\right)^{2}\leq\left(\frac{1}{x_{n-1}}\right)\left(\frac{1}{x_{n+1}}\right)$
Then the reciprocal sequence in Example 2 is a counterexample to Theorem 2(a) of Asai, Kubo, & Kuo [3] and to Proposition 2.1(b) of Došlić [12] and to Theorem 4.4 of Gajdzica, Miska, & Ulas [17].
Example 2. Define the sequence $\{b_{n}\}$ as the reciprocal of $a_{n}=2^{n+1}-1$ (as in Example 1) for $n=0,1,2,\ldots$ Then, because $a_{n}$ is strictly log-concave, the reciprocal sequence $b_{n}$ is strictly log-convex: $b_{n}^{2}>b_{n-1}b_{n+1}$. However, the sequence does not satisfy Condition (3):
$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}\qquad\iff\qquad(2^{n+1}-1)^{\frac{-1}{n+1}}>(2^{n+2}-1)^{\frac{-1}{n+2}}.$
Further, $b_{n}$ is not deficient (in fact, $b_{n}$ is abundant):
$(2^{n+1}-1)^{-1}(2^{m+1}-1)^{-1}>(2^{n+m}-1)^{-1}$
for $n,m\geq 1$.
Consider what happens to the conditions of Theorem 1.1 if the sequence $x_{n}$ is replaced with its reciprocal sequence. The analogue to Condition (2) for log-convexity is similarly sufficient to ensure that a sequence is deficient. Consider the following equivalent formulation of Theorem 1.1 in terms of log-convexity:
Theorem 1.2.
Let $\{y_{n}\}$ be a sequence of positive real numbers and let $N$ be a positive integer such that for all $n>N$, the sequence satisfies $y_{n}^{2}\leq y_{n-1}y_{n+1}$. Then for all $n>N$ and $m\geq 0$, the sequence $\{y_{n}\}$ satisfies:
$\left(y_{N}\right)^{\frac{m}{n-N+m}}\left(y_{n+m}\right)^{\frac{n-N}{n-N+m}}\geq y_{n}\geq y_{N}\left(\frac{y_{N+1}}{y_{N}}\right)^{n-N}.$
Moreover, if
$(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}},$ (3)
then $y_{n+m}\geq y_{n}y_{m}$ for all $n,m>N$.
Again, note that if $\{y_{n}\}$ is log-convex for all $n\geq 1$, then
$y_{1}^{\frac{m}{n+m-1}}(y_{n+m})^{\frac{n-1}{n+m-1}}\geq y_{n}\geq y_{1}\left(\frac{y_{2}}{y_{1}}\right)^{n-1}$
for all $n$ and $m$. Example 2 demonstrates that Condition (3) in Theorem 1.2, for a log-convex sequence to be deficient, is required. If a log-convex sequence does not satisfy $(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\geq x_{n}x_{m}$ for all $n,m\geq N+1$.
1.1. Sufficient but not Necessary
Theorem 1.1 establishes the condition of log-concavity together with Condition (2) as sufficient conditions that a sequence is abundant, but are they necessary conditions? In fact, there are sequences that are not log-concave (nor log-convex) yet are abundant.
Example 3. Define the sequence $a_{n}$ by
$a_{n}=\begin{cases}2,&\text{if}\ n\equiv 0\mod{4}\\ 3,&\text{if}\ n\equiv 1,3\mod{4}\\ 4,&\text{if}\ n\equiv 2\mod{4}\end{cases}$
$a_{n}=2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,\ldots\qquad\text{for}\ n=0,1,2,\ldots$
The inequality $a_{n}^{2}\geq a_{n-1}a_{n+1}$ fails whenever $n\equiv 0\mod{4}$, and so $\{a_{n}\}$ is never log-concave. Similarly, the inequality $a_{n}^{2}\leq a_{n-1}a_{n+1}$ fails whenever $n\equiv 2\mod{4}$, and so $\{a_{n}\}$ is never log-convex. However, the sequence $\{a_{n}\}$ satisfies the inequality $a_{n}a_{m}\geq a_{n+m}$ for all $n$ and $m$. Note that $(a_{n+1})^{\frac{1}{n+1}}\leq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 0,3\mod{4}$ and $(a_{n+1})^{\frac{1}{n+1}}\geq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 1,2\mod{4}$.
Analogously, Theorem 1.2 establishes the condition of log-convexity together with Condition (3) as sufficient conditions that a sequence is deficient, but are they necessary conditions? In fact, there are sequences that are neither log-concave nor log-convex yet are deficient.
Example 4. The Fibonacci sequence defined by $F_{0}=0$, $F_{1}=1$, and $F_{n}=F_{n-1}+F_{n-2}$, is log-concave for odd indices and log-convex for even indices (known as log-Fibonacci [3]). This is demonstrated in the literature by Cassini’s identity [38]:
$F_{n+1}F_{n-1}-F_{n}^{2}=(-1)^{n}$
for all $n$. The Fibonacci sequence is also known [22] to satisfy the identity
$F_{n+m}=F_{m+1}F_{n}+F_{m}F_{n-1}$
for all $n$ and $m$. Note that $F_{m+1}F_{n}+F_{m}F_{n-1}>F_{n}F_{m}$. Then, besides the single instance of equality at $(n,m)=(1,1)$, the Fibonacci sequence satisfies the inequality $F_{n+m}>F_{n}F_{m}$ for all $n$ and $m$. Note that for all $n>1$, the sequence $\{(F_{n})^{\frac{1}{n}}\}$ is a strictly increasing sequence that approaches the golden ratio in the limit, hence $(F_{n+1})^{\frac{1}{n+1}}<(F_{n+2})^{\frac{1}{n+2}}$.
1.2. Other Log-Concave Sequences
There are many examples of log-concave sequences that, in light of Theorem 1.1, are now known to be abundant. Log-concavity has been established for Stirling numbers of the first and second kind [34], the Hyperfibonacci and Hyperlucas numbers [39], the HyperPell and HyperPell-Lucas numbers [1], and the figurate and central figurate numbers [20, 19] to name a few. A more general class of sequences is proven log-concave by Sullivan [33] who considered sequences $\{\alpha(n)\}$ of the form $\alpha(n)=n^{a}\exp(cn^{b})$ for various values of $a,b$ and $c$.
It is natural to consider restricted partition functions, frequently denoted $p_{A}(n)$, where $A$ is some parameter limiting which values $\lambda$ can take. For instance, the integer $4$ could be partitioned among the restricted set of prime numbers, or perfect squares, or powers of $2$, or odd numbers, etc:
[…]
$=1^{2}+1^{2}+1^{2}+1^{2}$ $=2^{1}+2^{1}$ $=1+1+1+1$
$=2^{0}+2^{0}+2^{0}+2^{0}$
Proofs that particular restricted partition functions are log-concave or abundant have typically relied on knowing their asymptotic. The $k$-regular partition function was proven abundant in 2016 by Beckwith & Bessenrodt [4], and log-concavity was proven in 2019 by Craig & Pun [9]. The $k$-colored partition function was proven both log-concave and abundant in 2018 by Chern, Fu, & Tang [8]. Dawsey & Masri [27] proved the Andrews smallest parts function is log-concave and abundant in 2017. In 2022, the plane partition function was proven abundant by Heim, Neuhauser, & Tröger [21] who also conjectured log-concavity; this was later proven that same year by Ono, Pujahari, & Rolen [31].
There are restricted partition functions for which log-concavity has been established, but which have not been proven to be abundant. Proofs of log-concavity for such functions are given by Dong & Ji [11] for the distinct partition function (equivalent to the partitions into odd parts function), by Engel [13] for the overpartition function, by Iskander, Jain, & Talvola [23] for the fractional partition function, by Gajdzica [15] for the multiset partition function, and independently by Benfield, Roy, & Paul [5] and by Sullivan [32] for the $k$-th power partition function.
This paper proves that these restricted partition functions are abundant. Where possible for these restricted partition functions, in the spirit of the original publication by Bessenrodt & Ono [30], this paper also calculates the finitely many values where the strict abundant inequality fails or is equal.
2. Proof of Theorem 1.1 and 1.2
Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence that is log-concave for all $n>N$. For a fixed $n>N$, assume $x_{k}^{2}\geq x_{k-1}x_{k+1}$ where $k=N+1,N+2,\dots,n$. By log-concavity, consider first the two inequalities $x_{n}^{2}\geq x_{n-1}x_{n+1}$ and $x_{n-1}^{2}\geq x_{n}x_{n-2}$. After some rearrangement, it follows that
$x_{n+1}\leq\frac{(x_{n})^{2}}{x_{n-1}}\leq\frac{1}{x_{n-1}}\left(\frac{(x_{n-1})^{2}}{x_{n-2}}\right)^{2}=\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}.$ (4)
Again by log-concavity, $x_{n-2}^{2}\geq x_{n-1}x_{n-3}$, which can be substituted into (4), yielding
$\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}\leq\frac{1}{(x_{n-2})^{2}}\left(\frac{(x_{n-2})^{2}}{x_{n-3}}\right)^{3}=\frac{(x_{n-2})^{4}}{(x_{n-3})^{3}}.$
Iterating this process yields the following inequality:
$x_{n+1}\leq\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}\leq\frac{(x_{n-2})^{4}}{(x_{n-3})^{3}}\leq\frac{(x_{n-3})^{5}}{(x_{n-4})^{4}}\leq\ldots\leq\frac{(x_{N+2})^{n-N}}{(x_{N+1})^{n-N-1}}\leq x_{N}\left(\frac{x_{N+1}}{x_{N}}\right)^{n+1-N}.$ (5)
Hence, $x_{n}$ is bounded from above: $x_{n}\leq x_{N}\left(\frac{x_{N+1}}{x_{N}}\right)^{n-N}$ for all $n>N$ This gives the right-hand side of (1).
To prove the left-hand side of (1), note that the inequalities $(x_{N+1})^{2}\geq x_{N}x_{N+2}$ and $(x_{N+2})^{2}\geq x_{N+1}x_{N+3}$ follow from log-concavity. After some substitution, these inequalities can be written as
$(x_{N+2})^{2}\geq x_{N+1}x_{N+3}\geq\left((x_{N})^{\frac{1}{2}}(x_{N+2})^{\frac{1}{2}}\right)x_{N+3}\Rightarrow x_{N+2}\geq(x_{N})^{\frac{1}{3}}(x_{N+3})^{\frac{2}{3}}.$ (6)
Again by log-concavity, $(x_{N+3})^{2}\geq x_{N+2}x_{N+4}$ which after substitution into (6) can be written as
$(x_{N+3})^{2}$ $\geq x_{N+2}x_{N+4}\geq\left((x_{N})^{\frac{1}{3}}(x_{N+3})^{\frac{2}{3}}\right)x_{N+4}\Rightarrow x_{N+3}\geq(x_{N})^{\frac{1}{4}}(x_{N+4})^{\frac{3}{4}}.$
This process may continues recursively, and as $n$ gets large,
$x_{n}\geq\left(x_{N}\right)^{\frac{1}{n-N+1}}\left(x_{n+1}\right)^{\frac{n-N}{n-N+1}}.$ (7)
Increasing the index from $n$ to $n+1$ yields
$x_{n+1}\geq\left(x_{N}\right)^{\frac{1}{n-N+2}}\left(x_{n+2}\right)^{\frac{n-N+1}{n-N+2}}$
which can be substituted into (7) and rearranged:
$x_{n}\geq\left(x_{N}\right)^{\frac{1}{n-N+1}}\left(x_{n+1}\right)^{\frac{n-N}{n-N+1}}$ $\geq\left(x_{N}\right)^{\frac{1}{n-N+1}}\left(\left(x_{N}\right)^{\frac{1}{n-N+2}}\left(x_{n+2}\right)^{\frac{n-N+1}{n-N+2}}\right)^{\frac{n-N}{n-N+1}}$
$=(x_{N})^{\frac{1}{n-N+1}+\frac{n-N}{(n-N+1)(n-N+2)}}(x_{n+2})^{\frac{n-N}{n-N+2}}.$
Iterating this process up to some index $n+m$ for some $m\geq 0$ yields:
$x_{n}\geq(x_{N})^{\frac{1}{n-N+1}+\frac{n-N}{(n-N+1)(n-N+2)}+\dots+\frac{n-N}{(n-N+m-1)(n-N+m)}}(x_{n+m})^{\frac{n-N}{n-N+m}}.$ (8)
Note that inequality (8) holds for all $n>N$ and all $m\geq 0$, and that the exponent of $x_{N}$ in (8) is a telescoping series:
$\frac{1}{n-N+1}+\frac{n-N}{(n-N+1)(n-N+2)}+\dots+\frac{n-N}{(n-N+m-1)(n-N+m)}$
$=(n-N)\left[\left(\frac{1}{n-N}-\frac{1}{n-N+1}\right)+\left(\frac{1}{n-N+1}-\frac{1}{n-N+2}\right)+\ldots\right.$
$\hskip 216.81pt\left.+\left(\frac{1}{n-N+m-1}-\frac{1}{n-N+m}\right)\right]$
$=(n-N)\left(\frac{1}{n-N}-\frac{1}{n-N+m}\right)$
$=\frac{m}{n-N+m}$
Thus, inequality (8) is reduced to:
$x_{n}\geq(x_{N})^{\frac{m}{n-N+m}}(x_{n+m})^{\frac{n-N}{n-N+m}}$ (9)
for all $n>N$. This gives the left-hand side of (1).
To prove the right hand side of Theorem 1.1, it is necessary to establish relations for all entries in the sequence, in particular, for $x_{0},x_{1},\dots,x_{N}$. To achieve this, define the sequence $\{a_{n}\}$ recursively by $a_{n}=x_{n}$ for all $n\geq N+1$ and $1\leq a_{n-1}\leq\frac{(a_{n})^{2}}{a_{n+1}}$ for $n=1,2,\ldots,N+1$. In particular, one may chose $a_{n}=x_{n}$ for all $n\geq N+1$ and for $n\leq N$,
$a_{N}:=\frac{(a_{N+1})^{2}}{a_{N+2}}=\frac{(x_{N+1})^{2}}{x_{N+2}}$
$a_{N-1}:=\frac{(a_{N})^{2}}{a_{N+1}}=\frac{\left(\frac{(x_{N+1})^{2}}{x_{N+2}}\right)^{2}}{x_{N+1}}=\frac{(x_{N+1})^{3}}{(x_{N+2})^{2}}$
$\vdots$
$a_{N-N}=a_{0}:=\frac{(x_{N+1})^{N+2}}{(x_{N+2})^{N+1}}.$
Hence, $(a_{n+1})^{2}\geq a_{n}a_{n+2}$ for all $n\geq 0$. A similar process to that outlined in (5) yields
$a_{n+1}\leq\frac{(a_{n-1})^{3}}{(a_{n-2})^{2}}\leq\frac{(a_{n-2})^{4}}{(a_{n-3})^{3}}\leq\frac{(a_{n-3})^{5}}{(a_{n-4})^{4}}\leq\ldots\leq\frac{(a_{2})^{n}}{(a_{1})^{n-1}}=a_{1}\left(\frac{a_{2}}{a_{1}}\right)^{n}.$ (10)
Also, by the same process outlined between equations (6) and (9) one yields
$a_{n}\geq(a_{1})^{\frac{m}{n+m-1}}(a_{m+n})^{\frac{n-1}{n+m-1}}.$ (11)
Switching the indices $n$ and $m$ in (11) gives the inequality:
$a_{m}\geq(a_{1})^{\frac{n}{m+n-1}}(a_{m+n})^{\frac{m-1}{m+n-1}}.$ (12)
Suppose for contradiction that the sequence $\{a_{n}\}$ is log-concave but the sequence $\{a_{n}\}$ does not satisfies the inequality $a_{n+m}\leq a_{n}a_{m}$ for all $n,m\geq 0$. Then, for some $m\geq 0$ and $n\geq 0$, the inequality $a_{n+m}>a_{n}a_{m}$ holds. Hence, for one such pair of $m$ and $n$, and from (9) and (12) it follows that
$a_{n+m}$ $>a_{n}a_{m}$
$\geq\left((a_{1})^{\frac{m}{n+m-1}}(a_{n+m})^{\frac{n-1}{n+m-1}}\right)\left((a_{1})^{\frac{n}{m+n-1}}(a_{m+n})^{\frac{m-1}{m+n-1}}\right)$
$=(a_{1})^{\frac{n+m}{n+m-1}}(a_{n+m})^{\frac{n+m-2}{n+m-1}}.$
Hence
$(a_{n+m})^{\frac{1}{n+m-1}}>(a_{1})^{\frac{1}{n+m-1}+1}.$ (13)
Note by Condition (2) that $a_{0}\geq 1$. And by log-concavity, $a_{1}^{2}\geq a_{0}a_{2}\geq a_{2}$ implying that $\frac{a_{2}}{a_{1}}\leq a_{1}$. Then from (10) it follows that
$(a_{n+m})^{\frac{1}{n+m-1}}\leq\left[a_{1}\left(\frac{a_{2}}{a_{1}}\right)^{n+m-1}\right]^{\frac{1}{n+m-1}}$ $\leq\left(\frac{a_{2}}{a_{1}}\right)\left(a_{1}\right)^{\frac{1}{n+m-1}}$
$\leq(a_{1})^{\frac{1}{n+m-1}+1}.$ (14)
Hence
Both (13) and (14) cannot be simultaneously true, this gives a contradiction. Hence $a_{m+n}\leq a_{m}a_{n}$ for all $m,n\geq 0$ and in particular $m,n\geq N+1$ completing the proof of Theorem 1.1. ∎
Consider the consequences of starting this proof with the sequence $y_{n}=x_{n}^{-1}$. Then $y_{n}$ is log-convex for all $n>N$ and the same proof can be adapted where the only change is the reversal of every inequality sign, completing the proof of Theorem 1.2. ∎
3. Various Partition Functions
This section considers restricted partition functions $P_{A}(n)$ that are known to be log-concave. It follows from Theorem 1.1 that these sequences are also abundant, provided they also satisfy Condition (2). This section considers a few, well-studied sequences and partition functions and determines for which small indices abundance fails.
3.1. The Distinct Partition Function, Partitions into Odd Parts
The Distinct Partition Function $P_{d}(n)$ counts the number of ways to partition $n$ into distinct parts, where for all $i$ and $j$, $\lambda_{i}\neq\lambda_{j}$.
$P_{d}(n)$ $=1,1,1,2,2,3,4,5,6,8,10,12,15,18,22,27,32,38,46,54,64,76,89,104,122,142,165,$
[…]
Curiously, Euler showed that $P_{d}(n)$ is equivalent to the number of partitions of $n$ using only odd integers [14, 25, 28]. For example, $P_{d}(4)=2$ since there are only two ways to partition $4$ into distinct parts: $4=4=3+1$ (and only two ways to partition $4$ into odd parts: $4=3+1=1+1+1+1$). In 2023, Dong & Ji [11] proved that the sequence $\{P_{d}(n)\}$ is log-concave for all $n\geq 33$. Note that $P_{d}(n)$ satisfies Condition (2) for $n=33$: $448^{\frac{1}{33}}>512^{\frac{1}{34}}$. The result follows from Theorem 1.1 and numerical computation for small indices.
Theorem 3.1.
The distinct partition function satisfies the inequality $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$ where instances of equality and failure occur as given in Table 1.
Table 1. Equality and Failure for $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$
Equality at $(a,b)$ Failure at $(a,b)$
$(1,1),(3,1),$ $a=1,2$
$a=5,6,7,8\quad b=3,$ $(4,3),(5,5)$
$a=15,16,17\quad b=4,$ $44\leq a\leq 14\quad b=$ 4
$a=6,7,8\quad b=5$
3.2. The Overpartition Function
The overpartition function, $\overline{p}(n)$, counts the number of partitions of $n$ in which the last occurrence of a number can be distinguished. For example, $\overline{p}(4)=14$:
$\ 1+1+1+1,\quad 1+1+1+\overline{1},\quad 1+1+2,\quad 1+1+\overline{2},\quad 1+\overline{1}+2,$
$1+\overline{1}+\overline{2},\quad 2+2,\quad 2+\overline{2},\quad 1+3,\quad 1+\overline{3},\quad\overline{1}+3,\quad\overline{1}+\overline{3},\quad 4,\quad\overline{4}.$
[…]
In 2014, Engel [13] proved that $\{\overline{p}(n)\}$ function is log-concave for $n\geq 2$. Note that $\overline{p}$ satisfies Condition (2) for $n=2$: $\sqrt{4}\geq\sqrt[3]{8}$. The result follows from Theorem 1.1 and numerical computation for small indices.
Theorem 3.2.
The overpartition function satisfies the inequality $\overline{p}(a)\overline{p}(b)\geq\overline{p}(a+b)$ where instances of equality occur at $(1,1)$ and $(1,2)$.
The following subsections, 3.3, 3.4, and 3.5, have been proven to satisfy log-concavity for sufficiently large $n$, however, the smallest such index $N$ such that each sequence is log-concave for all $n\geq N$ is either unknown or only conjectured. On the other hand, it is known that the asymptotics of these sequences are each of the form $n^{a}e^{c\cdot n^{b}}$ for real numbers $a,b,$ and $c$ where $0<b\leq 1$. This implies that the $n$th root of these sequences are each decreasing, hence, there is an index $N$ such that $x_{N+1}^{\frac{1}{N+1}}\geq x_{N+2}^{\frac{1}{N+2}}$.
3.3. Fractional Partition Function
Let $\alpha$ be a real number and define the fractional partition function $p_{\alpha}(n)$ by
$\sum_{n=0}^{\infty}p_{\alpha}(n)q^{n}=(q;q)_{\infty}^{\alpha}$
In 2019 Huat & Chan [6] studied the properties of $p_{\alpha}(n)$ where alpha is a rational number. The next year, Iskander, Jain & Talvola [23] proved that for any real $\alpha$ there exists an integer $N_{\alpha}$ such that the fractional partition function $p_{\alpha}(n)$ is log-concave $n>N_{\alpha}$.
Theorem 3.3.
There is an index $M_{\alpha}$ such that the sequence $p_{\alpha}(n)$ satisfies the inequality $p_{\alpha}(n)p_{\alpha}(m)>p_{\alpha}(n+m)$ for all $n,m>M_{\alpha}$
3.4. The Multiset Partition Function
Let $k$ be a fixed positive integer and let $\mathcal{A}=\{a_{i}\}_{i=1}^{\infty}$ be a non-decreasing sequence of positive integers and define a restricted partition of $n$ the sequence of positive integers $\lambda_{1},\lambda_{2},\ldots,\lambda_{j}$ such that $n=\lambda_{1}+\lambda_{2}+\ldots+\lambda_{j}$ where each $\lambda_{i}$ belongs to the multiset $\{a_{1},a_{2},...,a_{k}\}$. The multiset partition function $p_{\mathcal{A}}(n,k)$ counts all partitions of $n$ among the first $k$ elements of the multiset $\mathcal{A}$. The sequence $\{p_{\mathcal{A}}(n)\}$ was studied in 2022 by Gajdzica [15] who proved log-concavity under certain conditions of $n$ and $k$. This result was futher improved by Gajdzica, Heim, & Markus [16].
Theorem 3.4.
There exists a sufficiently large index such that the multiset partition function satisfies the inequality $p_{\mathcal{A}}(n)p_{\mathcal{A}}(m)\geq p_{\mathcal{A}}(n+m)$ under the same conditions Gajdzica gives in his Theorem 1.1 [15].
3.5. The $k$-th Power Partition Function
The $k$-th power partition function $p^{k}(n)$ counts the number of partitions of $n$ among perfect $k$-th powers. For example, $p^{k}(n)=p(n)$ when $k=1$, and for $k=2$, $p^{2}(n)$ restricts $\lambda$ to perfect squares, etc.
$p^{1}(4)=5$ $p^{2}(4)=2$ $p^{k\geq 3}(4)=1$
$4$ $2^{2}$ $1^{k}+1^{k}+1^{k}+1^{k}$
$3+1$ $1^{2}+1^{2}+1^{2}+1^{2}$
$2+2$
$2+1+1$
$1+1+1+1$
Independently, both Benfield, Paul, & Roy [5] and Sullivan [32] proved that $p^{k}(n)$ is log-concave for all $k$ and sufficiently large $n$.
Theorem 3.5.
There exists an index $M_{k}$ such that for all $k$, the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ holds for $a,b>M_{k}+1$.
What is the smallest index for which log-concavity fails? Define for $p^{k}(n)$, the smallest index $N_{k}$ such that for all $n>N_{k}$, the sequence ${p^{k}(n)}$ is log-concave.
Conjecture 3.6.
The smallest $N_{k}$ such that $p^{k}(n)$ is log-concave for all $n>N_{k}$ for $k~{}\leq~{}6$ is given by:
Table 2. Smallest $N_{k}$ for $k=1,\ldots,6$, OEIS: A346160 [37]
$p^{k}(n)$ $p^{1}(n)$ $p^{2}(n)$ $p^{3}(n)$ $p^{4}(n)$ $p^{5}(n)$ $p^{6}(n)$
$N_{k}$ $25$ $1041$ $15655$ $637854$ $2507860$ $35577568$
Conjecture 3.7 gives instances $(a,b)$ where the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ is equal and when it fails for $k=1,2,3$. The case $k=1$ was proven by Bessenrodt & Ono [30]. Note that equality always holds for $a+b<2^{k}$.
Conjecture 3.7.
The $k$-th power partition function satisfies the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ with instances of failure or equality as given in Table 3:
Table 3. Equality and Failure for $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$
$p_{k}(n)$ Equality Failure
$a<2^{1},\ a+b>2^{1}$
$p_{1}(n)$ $(2,7),(2,6),(3,4)$ $a=2,3,4,5\quad b=2$
$a=3,5\quad b=3$
$a=5,6,7\quad b=4$ $a<2^{2},\ a+b>2^{2}$
$p_{2}(n)$ $a=5,6,8,11,15\quad b=5$ $(7,5)$
$a=8,10,12,14\quad b=6$ $a=6,7,11,15\quad b=6$
$a=8,9,12,13,19\quad b=7$ $a=7,10,11,14,15\quad b=7$
$a=15,23\quad b=9$ $a<2^{3},\ a+b>2^{3}$
$p_{3}(n)$ $a=14,15,22,23\quad b=10$ $a=15,23\quad b=12$
$a=13,14,15,21,22,23\quad b=11$ $a=14,15,22,23,b=13$
$a=12,13,14,20,21,22\quad b=12$ $a=14,15,21,22,23\quad b=14$
$a=13,19,20,21,23\quad b=13$ $a=15,20,21,22,23\quad b=15$
$a=18,19,20,26\quad b=14$
$a=17,18,19,26\quad b=15$
$(23,20),(22,21),(23,21),(22,22),(23,22),(23,23)$
4. Log-polynomial Sequences
In a more general setting, Sullivan [33] made interesting connections between the data of log-polynomial sequences and log-concavity/log-convexity. A sequence $\alpha(n)$ is log-polynomial of degree $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ for $\kappa=\pm 1$ if there exists $\delta(n)>0$, $A(n)$, and $g_{k}(n)$ for $k=3,4,\ldots,m$ such that
$\log{\left(\frac{\alpha(n+j)}{\alpha(n)}\right)}=A(n)j+\kappa\cdot\delta(n)^{2}j^{2}+\sum_{k=3}^{m}g_{k}(n)j^{k}+o\left(\delta(n)^{m+1}\right)$
for $j=1,2,\ldots,m+1$ where both $\delta(n)\rightarrow 0$ and $g_{k}(n)=o\left(\delta(n)^{k}\right)$ as $n\rightarrow\infty$. Sullivan [33] showed that if $\alpha(n)$ is log-polynomial for degrees $m\geq 2$, then $\alpha(n)$ is log-concave for large $n$ whenever $\kappa=-1$ and is log-convex for large $n$ whenever $\kappa=1$. In particular, Sullivan showed [33] that if $\alpha(n)=n^{a}e^{c\cdot n^{b}}$ for real numbers $a$, $b$, and $c$ where $a$ and $c$ are not both zero, $20<b<$ 2, and $b(b-1)c\neq 0$, then $\alpha(n)$ is log-polynomial with data $\{A(n),\kappa,\delta(n)\}$ where
$A(n)=-\frac{a}{n}+bc\cdot n^{b-1}\quad\ \quad\kappa=\text{sgn}(b(b-1)c)\quad\ \quad\kappa\cdot\delta(n)^{2}=-\frac{a}{2n^{2}}+\frac{b(b-1)c}{2n^{2-b}}$
for all degrees $m\geq 2$.
Theorem 4.1.
If a sequence $\alpha(n)$ is log-polynomial for degrees $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ and $0<b\leq 1$, then for sufficiently large integers $n$ and $m$, the sequence satisfies $\alpha(n)\alpha(m)\geq\alpha(n+m)$ whenever $\kappa=-1$ and satisfies $\alpha(n)\alpha(m)\leq\alpha(n+m)$ whenever $\kappa=1$ with finitely many instances of equality or failure.
5. Sufficient But Not Necessary Condition for Partition Functions
Although many restricted partition functions are log-concave (and hence abundant), there are restricted partition functions that are neither log-concave nor log-convex yet still appear to be abundant. Below is an empirical result for a particular restricted partition function.
5.1. The $m$-ary Partition Function
The $m$-ary partition function $b^{m}(n)$ counts the number of partitions of $n$ among powers of $m>1$. Asymptotics for the $m$-ary partitions functions were established by Kachi [24] in 2015. The most famous of the $m$-ary partition functions is the binary partition function which counts the partitions of $n$ among powers of $2$. For example, $b^{2}(4)=4$, $b^{3}(4)=2$, $b^{4}(4)=2$ and for $m\geq 5$, $b^{m}(4)=1$:
$b^{2}(4)=4$ $b^{3}(4)=2$ $b^{4}(4)=2$ $b^{m\geq 5}(4)=1$
$2^{2}$ $3^{1}+3^{0}$ $4^{1}$ $m^{0}+m^{0}+m^{0}+m^{0}$
$2^{1}+2^{1}$ $3^{0}+3^{0}+3^{0}+3^{0}$ $4^{0}+4^{0}+4^{0}+4^{0}$
$2^{1}+2^{0}+2^{0}$
$2^{0}+2^{0}+2^{0}+2^{0}$
In 2019, Ganguly [18] showed that $b^{m}(n)$ is neither log-concave nor log-convex. Yet, for each $m$, the sequence $\{b^{m}(n)\}_{n\in\mathbb{N}}$ appears to be abundant with only finitely many instances of equality or failure.
Conjecture 5.1.
For each $m$, the sequence given by $\{b^{m}(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $b^{m}(\alpha)b^{m}(\beta)~{}\geq~{}b^{m}(\alpha+\beta)$ with instances of equality or failure as given by Table 4.
Table 4. Equality and Failure for $b^{m}(\alpha)b^{m}(\beta)\geq b^{m}(\alpha+\beta)$
$b^{m}(n)$ Equality at $(a,b)$ Failure at $(\alpha,\beta)$
$b^{2}(n)$ $(2,2),(3,2),(3,9)$ $a,b\leq 1$,
$(3,3),(5,3),(7,3)$
$b^{3}(n)$ $(7,8),(8,8)$ $a,b\leq 2$,
$(4,5),(4,8),(5,5),(5,7),(5,8)$
$b^{4}(n)$ $(5,7),(5,11),(5,15),(6,6)$ $a,b\leq 3$,
$(6,7),(6,10),(6,11),(6,14)$
$(6,15),(7,7),(7,9),(7,10)$
$(7,11),(7,13),(7,14),(7,15)$
And for $m\geq 5$, Failure only occurs for $\alpha,\beta\leq m$. For $m=5$, equality occurs for pairs $(\alpha,\beta)$ where $12m-2\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$. For $m\geq 6$, equality occurs for pairs $(\alpha,\beta)$ where $12m-3\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$.
For example,
$b^{5}(n)\ \text{has equality at}$ $b^{6}(n)\ \text{has equality at}$
$(8,7),(8,8)$ $(9,9)$
$(9,6),(9,7),(9,8),(9,9)$ $(10,8),(10,9),(10,10)$
$(11,7),(11,8),(11,9),(11,10),(11,11)$
6. Final Thoughts
Log-concavity (or log-convexity) along with Condition (2) (or with Condition (3)) have been shown to be sufficient conditions for a sequence to be abundant (or deficient). Perhaps more interesting are sequences that are neither log-concave nor log-convex yet are still abundant or deficient. There are essentially seven types of sequences that are neither log-concave nor log-convex, according to whether the sequence adheres to Conditions (2) or (3) or neither, and whether it is abundant, deficient, or neither.
The first two types of sequences are those that are neither log-concave nor log-convex and satisfy neither Conditions (2) nor (3) yet are abundant or deficient. Example 3 presents a sequence that is neither log-concave nor log-convex, and satisfies neither Conditions (2) nor (3), yet, the sequence is deficient for all indices $n>1$. Taking the reciprocal sequence of Example 3 reverses the inequalities in each property, producing a sequence that is neither log-concave nor log-convex, and satisfies neither Conditions (2) nor (3), yet, the sequence is abundant for all indices $n>1$.
The next two types of sequences are those that are neither log-concave nor log-convex, but do satisfy Condition (2), and are abundant or deficient. Taking the reciprocal the sequence found in Example 4 (the Fibonacci sequence) reverses the inequalities in each property, producing a sequence that is neither log-concave nor log-convex, but does satisfy Condition (2), and is abundant.
Question 1.
Does there exist a sequence that is neither log-concave nor log-convex, yet satisfies Condition (2) and is deficient?
The next two types of sequences are those that are neither log-concave nor log-convex, but do satisfy Condition (3), and are abundant or deficient. Example 4, the Fibonacci sequence, is neither log-concave nor log-convex, but does satisfy Condition (3), and is deficient.
Question 2.
Does there exist a sequence that is neither log-concave nor log-convex, yet satisfies Condition (3) and is abundant?
The last type of sequences are those that are neither log-concave nor log-convex, satisfy neither Conditions (2) nor (3), and are neither abundant nor deficient. There are certainly such sequences; more investigation is required to categorize exactly which sequences satisfy these curious conditions.
Data Sharing and Conflicts of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest. Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Acknowledgements
The authors would like to extend sincere gratitude to the anonymous referee, whose keen observations and criticism throughout the review process has greatly improved the quality of this paper.

### v4

@@
 Abstract.
-The partition function $p(n)$ and many of its related restricted partition functions have recently been shown independently to satisfy log-concavity: $p(n)^{2}\geq p(n-1)p(n+1)$, and satisfy the inequality: $p(n)p(m)\geq p(n+m)$ with only finitely many instances of equality or failure.
+The partition function $p(n)$ and many of its related restricted partition functions have recently been shown independently to satisfy log-concavity: $p(n)^{2}\geq p(n-1)p(n+1)$ for $n\geq 26$, and satisfy the inequality: $p(n)p(m)\geq p(n+m)$ for $n\geq m\geq 2$ with only finitely many instances of equality or failure.
 This paper proves that this is no coincidence, that any log-concave sequence $\{x_{n}\}$ satisfying a particular initial condition likewise satisfies the inequality $x_{n}x_{m}\geq x_{n+m}$.
@@
 Recent interest regarding the various properties of $p(n)$ led to a conjecture by Chen [9], and subsequent proof by DeSalvo & Pak [12], that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ is log-concave for $n>25$.
-A sequence of integers $\{x_{n}\}_{n\in\mathbb{N}}$ is log-concave if $x_{n}^{2}~{}\geq~{}x_{n-1}x_{n+1}$ and is is log-convex if $x_{n}^{2}\leq x_{n-1}x_{n+1}$ for all $n$.
+A sequence of integers $\{x_{n}\}_{n\in\mathbb{N}}$ is log-concave if $x_{n}^{2}~{}\geq~{}x_{n-1}x_{n+1}$ and is log-convex if $x_{n}^{2}\leq x_{n-1}x_{n+1}$ for all $n\geq n_{0}$ for some $n_{0}\in\mathbb{N}$.
 Log-concavity had been previously established in 1978 by Nicholas [32] for the partition function, but there was a revitalization of interest following the result of DeSalvo & Pak [12].
-New results found in the work of Bessenrodt & Ono [30] in 2014 proved that the sequence $\{p(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $p(n)p(m)\geq p(n+m)$ where instances of equality occur at $(2,6),(3,4)$, and instances of failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(5,5)$ and nowhere else.
-For convenience, this paper uses the notation: a sequence $\{x_{n}\}$ is multiplicatively abundant (or simply abundant) if $x_{n}x_{m}\geq x_{n+m}$, and is multiplicatively deficient (or simply deficient) if $x_{n}x_{m}\leq x_{n+m}$ for all but finitely many $n$ and $m$.
-In these terms, Bessenrodt & Ono [30] proved that $\{p(n)\}$ is abundant.
-Proofs that the partition function is log-concave or abundant have so far relied on an analytic result of Rademacher type due to Lehmer [26] until 2017 when A.
+New results found in the work of Bessenrodt & Ono [6] in 2014 proved that the sequence$\{p(n)\}_{n\in\mathbb{N}}$ satisfies the inequality
+$p(n)p(m)\geq p(n+m)\quad\text{for all}\quad(m,n)\quad\text{with}\quad n\geq m\geq 2$ (1)
+where instances of equality occur at $(2,6),(2,7),(3,4)$, and instances of failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(3,5)$ and nowhere else.
+For convenience, this paper uses the notation: a sequence $\{x_{n}\}$ is multiplicatively abundant (or simply abundant) if $x_{n}x_{m}\geq x_{n+m}$, and is multiplicatively deficient (or simply deficient) if $x_{n}x_{m}\leq x_{n+m}$ for all $n\geq m\geq n_{0}$ but finitely many $n\geq m\geq n_{0}$.
+In these terms, Bessenrodt & Ono [6] proved that $\{p(n)\}$ is abundant for all $n\geq m\geq 2$.
+Proofs that the partition function is log-concave or abundant initially relied on an analytic result of Rademacher type due to Lehmer [28] until 2017 when A.
 Alanazi, S.M.
@@
 Many sequences (especially sequences arising from restricted partitions) are sporadic for small indices but become and stay log-concave for sufficiently large indices.
-The main goal of this paper is to prove that any sequence that is eventually log-concave is also eventually abundant, provided the sequence also satisfies a necessary (but not sufficient) initial condition.
+The interaction between log-behavior and multiplicative behavior in the setting of general sequences was studied in the early 2000’s by Asai, Kubo, & Kuo [3] who collectively established inequalities for Bell numbers, Motzkin numbers, Fine numbers, Fresnel numbers, Apèry numbers, large Schröder numbers, central Delannoy numbers, and Baxter permutations.
+Recently, Gajdzica, Miska, & Ulas [18] studied similar interactions in a more general setting.
+The main goal of this paper is to prove that any sequence that is eventually log-concave is also eventually abundant, provided the sequence also satisfies a sufficient initial conditions.
 Theorem 1.1.
-Let $\{x_{n}\}$ be a sequence of positive real numbers and let $N$ be a positive integer such that for all $n>N$, the sequence satisfies $x_{n}^{2}\geq x_{n-1}x_{n+1}$.
+Let $\{x_{n}\}$ be a sequence of positive real numbers and let $N$ be a non-negative integer such that for all $n>N$, the sequence satisfies $x_{n}^{2}\geq x_{n-1}x_{n+1}$.
 Then for all $n>N$ and $m\geq 0$, the sequence $\{x_{n}\}$ satisfies:
 $(x_{N})^{\frac{m}{n-N+m}}(x_{n+m})^{\frac{n-N}{n-N+m}}\leq x_{n}\leq x_{N}\left(\frac{x_{N+1}}{x_{N}}\right)^{n-N}.$ (2)
-Moreover, if
-$(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$ (2)
-then $x_{n+m}\leq x_{n}x_{m}$ for all $n,m>N$.
-Note that if $\{x_{n}\}$ is log-concave for all $n\geq 1$, then
-$x_{1}^{\frac{m}{n+m-1}}(x_{n+m})^{\frac{n-1}{n+m-1}}\leq x_{n}\leq x_{1}\left(\frac{x_{2}}{x_{1}}\right)^{n-1}$
-for all non-negative integer $m$.
-In light of Theorem 1.1 and because log-concavity has been established for $n>25$ [10], to show the partition function is abundant it suffices to show that Condition (2) holds for $p(26)=2436$ and $p(27)=3010$: it is easy to verify that in fact, $2436^{\frac{1}{26}}>3010^{\frac{1}{27}}$.
-An example demonstrates that Condition (2) in Theorem 1.1, for a log-concave sequence to be abundant, is required.
-If a log-concave sequence does not satisfy $(x_{N+1})^{\frac{1}{N+1}}\geq(x_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\leq x_{n}x_{m}$ for all $n,m\geq N+1$.
-Example 1. Define the sequence $\{a_{n}\}$ by $a_{n}=2^{n+1}-1$ for $n=0,1,2,\ldots$
-$a_{n}=1,3,7,15,31,63,127,255,511,1023,2047,4095,8191,16383,32767,65535,\ldots$
-The sequence is strictly log-concave:
-$12^{2n+2}-2^{n+2}+1=(2^{n+1}-1)^{2}>(2^{n}-1)(2^{n+2}-1)=2^{2n+2}-2^{n}-2^{n+2}+$ 1
-However, the sequence $\{(2^{n+1}-1)^{\frac{1}{n+1}}\}$ for $n=0,1,2,\ldots$, is a strictly increasing sequence whose limit equals $2$ as $n$ tends to infinity.
-Hence, $a(n)$ does not satisfy Condition (2):
-$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}.$
+Moreover, if there exists an integer $k\geq 0$ such that
+$(x_{N+k})^{N+k+1}\geq(x_{N+k+1})^{N+k}$ (3)
+then $x_{n+m}\leq x_{n}x_{m}$ for all $n,m\geq N+k$.
+In addition, $\{x_{n}\}$ is abundant for all $m,n\geq d$ with at most finite number of failures $(m,n)\in[d,M+1]\times[d,N+k]\bigcup[d,N+k]\times[d,M+1]$, if $d\leq N+k$ is the smallest positive integer such that
+$\left(\frac{x_{M+1}}{x_{M}}\right)^{m}\leq x_{m}$ (4)
+for all $d\leq m\leq N+k$ and for some $M\geq N+k$.
+Remarks: 1) If a log-concave sequence does not satisfy condition (3), i.e., $(x_{N+k})^{N+k+1}\geq(x_{N+k+1})^{N+k}$ for all $k\geq 0$, then it is not true that $x_{n+m}\leq x_{n}x_{m}$.
+For example, define the sequence $\{a_{n}\}$ by $a_{n}=2^{n}-1/2$ for $n=0,1,2,\ldots$ The sequence is strictly log-concave:
+$\left(2^{n}-\frac{1}{2}\right)^{2}=2^{2n}-2^{n}+\frac{1}{4}>2^{2n}-2^{n}-2^{n-2}+\frac{1}{4}=\left(2^{n+1}-\frac{1}{2}\right)\left(2^{n-1}-\frac{1}{2}\right).$
+However, the sequence $\{(2^{n}-1/2)^{\frac{1}{n}}\}$ for $n=1,2,3,\ldots$, is a strictly increasing sequence.
+Hence, $a(n)$ does not satisfy Condition (3).
+Indeed, $(2^{n+k}-1/2)^{n+k+1}<(2^{n+k+1}-1/2)^{n+k}$ for all $n,k\geq 0$.
 In this particular case, $\{a_{n}\}$ is also not abundant (in fact, $\{a_{n}\}$ is deficient):
-$(2^{n+1}-1)(2^{m+1}-1)=2^{n+m+2}-2^{n}-2^{m}+1<2^{n+m}-1$
-for $n,m\geq 1$.
-The interaction between log-behavior and multiplicative behavior in the setting of general sequences was been studied in the early 2000’s by Asai, Kubo, & Kuo [3] and by Došlić [12] who collectively established inequalities for Bell numbers, Motzkin numbers, Fine numbers, Fresnel numbers, Apèry numbers, large Schröder numbers, central Delannoy numbers, and Baxter permutations.
-Both Asai, Kubo, & Kuo [3] and Došlić [12] give a proof that any log-convex sequence is deficient, provided $x_{0}=1$.
-More recently, a similar version of this proof emerged in 2023 by Gajdzica, Miska, & Ulas [17] with the slight modification that $x_{0}\geq 1$.
-Unfortunately, there is some misunderstanding that every log-convex sequence is deficient.
-Note that the reciprocal of a log-concave sequence of positive numbers is a log-convex sequence, and vice versa.
-$x_{n}^{2}\geq x_{n-1}x_{n+1}\qquad\iff\qquad\left(\frac{1}{x_{n}}\right)^{2}\leq\left(\frac{1}{x_{n-1}}\right)\left(\frac{1}{x_{n+1}}\right)$
-Then the reciprocal sequence in Example 2 is a counterexample to Theorem 2(a) of Asai, Kubo, & Kuo [3] and to Proposition 2.1(b) of Došlić [12] and to Theorem 4.4 of Gajdzica, Miska, & Ulas [17].
-Example 2.
-Define the sequence $\{b_{n}\}$ as the reciprocal of $a_{n}=2^{n+1}-1$ (as in Example 1) for $n=0,1,2,\ldots$ Then, because $a_{n}$ is strictly log-concave, the reciprocal sequence $b_{n}$ is strictly log-convex: $b_{n}^{2}>b_{n-1}b_{n+1}$.
-However, the sequence does not satisfy Condition (3):
-$(2^{n+1}-1)^{\frac{1}{n+1}}<(2^{n+2}-1)^{\frac{1}{n+2}}\qquad\iff\qquad(2^{n+1}-1)^{\frac{-1}{n+1}}>(2^{n+2}-1)^{\frac{-1}{n+2}}.$
-Further, $b_{n}$ is not deficient (in fact, $b_{n}$ is abundant):
-$(2^{n+1}-1)^{-1}(2^{m+1}-1)^{-1}>(2^{n+m}-1)^{-1}$
-for $n,m\geq 1$.
+$(2^{n}-1/2)(2^{m}-1/2)=2^{n+m}-2^{n-1}-2^{m-1}+\frac{1}{4}<2^{n+m}-1/2$
+for $n\geq m\geq 0$.
+2) Suppose $N=0$, i.e., $\{x_{n}\}$ is log-concave for all $n\geq 1$, then
+$x_{0}^{\frac{m}{n+m}}(x_{n+m})^{\frac{n}{n+m}}\leq x_{n}\leq x_{0}\left(\frac{x_{1}}{x_{0}}\right)^{n}$
+for all non-negative integers $m$.
+This give an upper bound for a log-concave sequence $\{x_{n}\}$.
+For example, a sequence with super exponential growth can not be log-concave.
+Moreover, $x_{n+m}\leq x_{n}x_{m}$ for all $n\geq m\geq k$ if $(x_{k})^{k+1}\geq(x_{k+1})^{k}$ for any fixed non-negative integer $k$.
+Specifically, if $k=0$, meaning that $x_{0}\geq 1$, then for all $n\geq m\geq 0$, $x_{n+m}\leq x_{n}x_{m}$.
+In 2001, Asai, Kubo, & Kuo [3, Theorem 2] [^1] While $x_{0}=1$ is mentioned in the theorem’s statement, the proof works for $x_{0}\leq 1$.
+To obtain the proof for the log-concave sequence, one must take the reciprocal of the log-convex sequence in [3].
+proved the case $N=0$ and $k=0$ of Theorem 1.1.
+Gajdzica, Miska, & Ulas [18, Theorem 4.4] proved this case again in $2023$ using a similar technique used in [3].
+It can be seen that Theorem 1.1 not only generalized the results of [3, Theorem 2] or [18, Theorem 4.4], it is also more effective.
+For instance, Bessenrodt-Ono result (1) can be obtained from Theorem 1.1 but it can not be derived from the $N=k=0$ case of Theorem 1.1.
+Instead one may obtain a weaker result such as $\{p(n+26)\}_{n\geq 0}$ is abundant (see [18, Example 4.5]).
+A weaker version of Theorem 1.1 for $k=0$ and $M=N$ was recently derived by Heim and Neuhauser [22].
+As we shall demonstrate in section 1.1, the result given in [22] does not imply that the log-concave sequences in example 5 are abundant for smaller indices.
+1.1. Some Examples of Theorem 1.1
+1) In light of Theorem 1.1 and because log-concavity has been established for $n>25$ in [12, 32], to show the partition function is abundant it suffices to show that Condition (3) holds for $p(25)=1958$ and $p(26)=2436$ with $N=25$ and $k=0$: it is easy to verify that in fact, $1958^{\frac{1}{25}}>2436^{\frac{1}{26}}$.
+It can be noted that this result only proves the partition function is abundant for $n\geq m\geq 25$ but not the result of Bessenrodt-Ono given in (1).
+To show (1) holds one needs to check condition (4), in particular, for $M=25$: $p(26)/p(25)\leq(p(m))^{1/m}$ for all $2\leq m\leq 25$.
+This can be easily verified.
+Next, by checking in the region $[2,26]\times[2,26]$, we see failure occur at $(2,2)$, $(2,3)$, $(2,4)$, $(2,5)$, $(3,3)$, $(3,5)$, $(3,2)$, $(4,2)$, $(5,2)$, $(5,3)$ and equality occur at $(2,6),(2,7),(3,4),(6,2),(7,2),(4,3)$.
+This gives the result of Bessenrodt-Ono (1).
+2) The overpartition function, $\overline{p}(n)$, counts the number of partitions of $n$ in which the last occurrence of a number can be distinguished. For example, $\overline{p}(4)=14$:
+$\ 1+1+1+1,\quad 1+1+1+\overline{1},\quad 1+1+2,\quad 1+1+\overline{2},\quad 1+\overline{1}+2,$
+$1+\overline{1}+\overline{2},\quad 2+2,\quad 2+\overline{2},\quad 1+3,\quad 1+\overline{3},\quad\overline{1}+3,\quad\overline{1}+\overline{3},\quad 4,\quad\overline{4}.$
+[…]
+Combining the result of Engel [14] and the fact $\overline{p}^{2}(1)\geq\overline{p}(0)\overline{p}(2)$ we see that $\{\overline{p}(n)\}$ is log-concave for $n>0$.
+Note that $\overline{p}$ satisfies both Conditions (3) and (4) for $N=k=0$.
+Then Theorem 1.1 give us that the overpartition function satisfies the inequality
+$\overline{p}(m)\overline{p}(n)\geq\overline{p}(m+n)\quad\text{forall}\quad m,n\geq 0$ (5)
+where instances of equality occur at $(1,1)$, $(1,2)$, $(2,1)$, $(m,0)$ and $(0,n)$. Note that (5) was proven using the asymptotic of the overpartition function by Liu & Zhang [29] in 2021.
+3) The distinct Partition Function $P_{d}(n)$ counts the number of ways to partition $n$ into distinct parts, where for all $i$ and $j$, $\lambda_{i}\neq\lambda_{j}$.
+$P_{d}(n)$ $=1,1,1,2,2,3,4,5,6,8,10,12,15,18,22,27,32,38,46,54,64,76,89,104,122,142,165,$
+[…]
+Curiously, Euler showed that $P_{d}(n)$ is equivalent to the number of partitions of $n$ using only odd integers [15, 27, 31].
+For example, $P_{d}(4)=2$ since there are only two ways to partition $4$ into distinct parts: $4=4=3+1$ (and only two ways to partition $4$ into odd parts: $4=3+1=1+1+1+1$).
+In 2023, Dong & Ji [13] proved that the sequence $\{P_{d}(n)\}$ is log-concave for all $n>32$.
+Note that $P_{d}(n)$ satisfies Condition (3) for $N=32$ and $k=0$: $390^{\frac{1}{32}}>448^{\frac{1}{33}}$ and condition (4) for $M=N=32$: $448/390\leq p_{d}(m)^{1/m}$ for all $3\leq m\leq 32$.
+Then from Theorem 1.1 and numerical computation gives the distinct partition function satisfies the inequality
+$P_{d}(m)P_{d}(n)\geq P_{d}(m+n)\quad\text{for all}\quad n\geq n>2$ (6)
+where instances of equality and failure occur as given in Table 1.
+Table 1. Equality and Failure for $P_{d}(m)P_{d}(n)\geq P_{d}(m+n)$
+Equality at $(m,n)$ Failure at $(m,n)$
+$(1,n)$; $n=1,3$ $(1,n),(2,n)$ ; $n\geq 2$
+$(3,n)$; $n=3,5,6,7,8$ $(1,2)$, $(3,4)$, $(5,5)$
+$(4,n)$; $n=15,16,17$ $(4,n)$; $4\leq n\leq 14,$
+$(5,n)$; $n=6,7,8$
+The inequality (6) together with the table 1 was proven by Beckwith & Bessenrodt in [4] using a Rademacher type asymptotic formula of $k$-regular partition.
+4) This is an example of $k>0$ case.
+Define the sequences $a_{0}=\frac{1}{2}$ and $a_{n}=2^{\sqrt{n}}$ for $n\geq 1$.
+Keep in mind that $a_{1}^{2}=2^{2}>\frac{1}{2}\times 2^{\sqrt{2}}=a_{0}a_{2}$, and that $a_{n+1}^{2}=2^{2\sqrt{n+1}}>2^{\sqrt{n}}\times 2^{\sqrt{n+2}}=a_{n}a_{n+2}$ for all $n\geq 1$ and that $a_{n+m}=2^{\sqrt{n+m}}<2^{\sqrt{n}}\times 2^{\sqrt{m}}=a_{n}a_{m}$ for all $n,m\geq 1$.
+In other words, $\{a_{n}\}$ fulfills the inequality $a_{n+m}<a_{n}a_{m}$ for all $n,m\geq 1$, but is a log-concave sequence for any $n$ with $a_{0}<1$.
+Since $a_{0}<1$, the result in [18, Theorem 4.4] is not applicable here.
+It is hopefully clear that $(a_{1})^{2}=2^{2}>2^{\sqrt{2}}=(a_{2})^{1}$ which satisfy Condition (3) in Theorem 1.1.
+5) Lastly, we give an example when $M>N+k$.
+Consider the sequence $a_{n}=\left(\frac{31}{25}\right)^{n}+\frac{1}{10^{6}}$ when $0\leq n\leq 24$ and $a_{n}=p(n)$ when $n\geq 25$.
+Here $p(n)$ is the partition function.
+It can be checked that $a_{n}^{2}\geq a_{n-1}a_{n-2}$ for $n\geq 25$.
+Also, it can be verified that $(a_{n})^{1/n}<\frac{a_{26}}{a_{25}}<\frac{a_{25}}{a_{24}}$ for all $1\leq n\leq 24$ but $(a_{n})^{1/n}>\frac{a_{27}}{a_{26}}$ for all $1\leq n\leq 25$.
+Thus, using Theorem 1.1, we can deduce that, for every $m,n\geq 1$, $a_{n+m}\leq a_{n}a_{m}$, with the exception of finitely many failures in $[1,25]\times[1,25]$.
+Note that, the result in [22] yields a weaker result, which only gives $a_{n+m}\leq a_{n}a_{m}$ for all $m,n\geq 24$.
 Consider what happens to the conditions of Theorem 1.1 if the sequence $x_{n}$ is replaced with its reciprocal sequence.
@@
 $\left(y_{N}\right)^{\frac{m}{n-N+m}}\left(y_{n+m}\right)^{\frac{n-N}{n-N+m}}\geq y_{n}\geq y_{N}\left(\frac{y_{N+1}}{y_{N}}\right)^{n-N}.$
-Moreover, if
-$(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}},$ (3)
-then $y_{n+m}\geq y_{n}y_{m}$ for all $n,m>N$.
-Again, note that if $\{y_{n}\}$ is log-convex for all $n\geq 1$, then
-$y_{1}^{\frac{m}{n+m-1}}(y_{n+m})^{\frac{n-1}{n+m-1}}\geq y_{n}\geq y_{1}\left(\frac{y_{2}}{y_{1}}\right)^{n-1}$
-for all $n$ and $m$.
-Example 2 demonstrates that Condition (3) in Theorem 1.2, for a log-convex sequence to be deficient, is required.
-If a log-convex sequence does not satisfy $(y_{N+1})^{\frac{1}{N+1}}\leq(y_{N+2})^{\frac{1}{N+2}}$, then it is not necessarily true that $x_{n+m}\geq x_{n}x_{m}$ for all $n,m\geq N+1$.
+Moreover, if there exists an integer $k\geq 0$ such that
+$(y_{N+k})^{N+k+1}\leq(y_{N+k+1})^{N+k},$ (7)
+then $y_{n+m}\geq y_{n}y_{m}$ for all $n,m\geq N+k$.
+In addition, $\{y_{n}\}$ is deficient for all $m,n\geq d$ with at most finite number of failures $(m,n)\in[d,M+1]\times[d,N+k]\bigcup[d,N+k]\times[d,M+1]$, if $d\leq N+k$ is the smallest positive integer such that
+$\left(\frac{y_{M+1}}{y_{M}}\right)^{m}\geq y_{m}$ (8)
+for all $d\leq m\leq N+k$ and for some $M\geq N+k$.
 1.2. Sufficient but not Necessary
@@
 Example 1. Define the sequence $a_{n}$ by
-$a_{n}=\begin{cases}2,&\text{if}\ n\equiv 0\mod{4}\\ 3,&\text{if}\ n\equiv 1,3\mod{4}\\ 4,&\text{if}\ n\equiv 2\mod{4}\end{cases}$
-$a_{n}=2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,2,3,4,3,\ldots\qquad\text{for}\ n=0,1,2,\ldots$
+$a_{n}=\begin{cases}2,&\text{if}\ n\equiv 0\mod{4}\\ 3,&\text{if}\ n\equiv 1,3\mod{4}\\ 4,&\text{if}\ n\equiv 2\mod{4}.\end{cases}$
 The inequality $a_{n}^{2}\geq a_{n-1}a_{n+1}$ fails whenever $n\equiv 0\mod{4}$, and so $\{a_{n}\}$ is never log-concave.
@@
 However, the sequence $\{a_{n}\}$ satisfies the inequality $a_{n}a_{m}\geq a_{n+m}$ for all $n$ and $m$.
-Note that $(a_{n+1})^{\frac{1}{n+1}}\leq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 0,3\mod{4}$ and $(a_{n+1})^{\frac{1}{n+1}}\geq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 1,2\mod{4}$.
+For $n>0$, note that $(a_{n+1})^{\frac{1}{n+1}}\leq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 0,3\mod{4}$ and $(a_{n+1})^{\frac{1}{n+1}}\geq(a_{n+2})^{\frac{1}{n+2}}$ whenever $n\equiv 1,2\mod{4}$.
 Analogously, Theorem 1.2 establishes the condition of log-convexity together with Condition (7) as sufficient conditions that a sequence is deficient, but are they necessary conditions?
@@
 Log-concavity has been established for Stirling numbers of the first and second kind [37], the Hyperfibonacci and Hyperlucas numbers [42], the HyperPell and HyperPell-Lucas numbers [1], and the figurate and central figurate numbers [21, 20] to name a few.
-A more general class of sequences is proven log-concave by Sullivan [33] who considered sequences $\{\alpha(n)\}$ of the form $\alpha(n)=n^{a}\exp(cn^{b})$ for various values of $a,b$ and $c$.
+A more general class of sequences is proven log-concave by O’Sullivan [36] who considered sequences $\{a(n)\}$ of the form $a(n)=n^{r}\exp(tn^{s})$ for various values of $r,s$ and $t$.
 It is natural to consider restricted partition functions, frequently denoted $p_{A}(n)$, where $A$ is some parameter limiting which values $\lambda$ can take.
@@
 $=2^{0}+2^{0}+2^{0}+2^{0}$
-Proofs that particular restricted partition functions are log-concave or abundant have typically relied on knowing their asymptotic.
-The $k$-regular partition function was proven abundant in 2016 by Beckwith & Bessenrodt [4], and log-concavity was proven in 2019 by Craig & Pun [9].
-The $k$-colored partition function was proven both log-concave and abundant in 2018 by Chern, Fu, & Tang [8].
+Proofs that particular restricted partition functions are log-concave or abundant or both have typically relied on knowing their asymptotic.
+The $k$-regular partition function for $2\leq k\leq 6$ was proven abundant in 2016 by Beckwith & Bessenrodt [4], and log-concavity was proven in 2019 by Craig & Pun [11] for large $n$.
+Proofs of log-concavity for the $k$-regular partition with $k=2$ and $3$ are given by Dong & Ji [13].
+The $k$-colored partition function was proven abundant in 2018 by Chern, Fu, & Tang [10].
+Log-concavity of the $k$-colored partition was proved by Bringmann, Kane, Rolen, & Tripp [7] which was conjectured in [10].
 Dawsey & Masri [30] proved the Andrews smallest parts function is log-concave and abundant in 2017.
@@
 There are restricted partition functions for which log-concavity has been established, but which have not been proven to be abundant.
-Proofs of log-concavity for such functions are given by Dong & Ji [11] for the distinct partition function (equivalent to the partitions into odd parts function), by Engel [13] for the overpartition function, by Iskander, Jain, & Talvola [23] for the fractional partition function, by Gajdzica [15] for the multiset partition function, and independently by Benfield, Roy, & Paul [5] and by Sullivan [32] for the $k$-th power partition function.
+Proofs of log-concavity for such functions are given by Iskander, Jain, & Talvola [25] for the fractional partition function, by Gajdzica [16] for the multiset partition function, and independently by Benfield, Roy, & Paul [5] and by O’Sullivan [35] for the $k$-th power partition function.
 This paper proves that these restricted partition functions are abundant.
-Where possible for these restricted partition functions, in the spirit of the original publication by Bessenrodt & Ono [30], this paper also calculates the finitely many values where the strict abundant inequality fails or is equal.
-2. Proof of Theorem 1.1 and 1.2
-Let $\{x_{n}\}_{n\in\mathbb{N}}$ be a sequence that is log-concave for all $n>N$.
-For a fixed $n>N$, assume $x_{k}^{2}\geq x_{k-1}x_{k+1}$ where $k=N+1,N+2,\dots,n$.
-By log-concavity, consider first the two inequalities $x_{n}^{2}\geq x_{n-1}x_{n+1}$ and $x_{n-1}^{2}\geq x_{n}x_{n-2}$.
+Where possible for these restricted partition functions, in the spirit of the original publication by Bessenrodt and Ono [6], this paper also calculates the finitely many values where the strict abundant inequality fails or is equal.
+2. Proof of Theorem 1.1
+Let us assume $x_{k}^{2}\geq x_{k-1}x_{k+1}$ where $k=N+1,N+2,\dots,n$.
+Now consider the two inequalities $x_{n}^{2}\geq x_{n-1}x_{n+1}$ and $x_{n-1}^{2}\geq x_{n}x_{n-2}$.
 After some rearrangement, it follows that
 $x_{n+1}\leq\frac{(x_{n})^{2}}{x_{n-1}}\leq\frac{1}{x_{n-1}}\left(\frac{(x_{n-1})^{2}}{x_{n-2}}\right)^{2}=\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}.$ (9)
-Again by log-concavity, $x_{n-2}^{2}\geq x_{n-1}x_{n-3}$, which can be substituted into (4), yielding
+Since $x_{n-2}^{2}\geq x_{n-1}x_{n-3}$, substituting this into (9), yielding
 $\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}\leq\frac{1}{(x_{n-2})^{2}}\left(\frac{(x_{n-2})^{2}}{x_{n-3}}\right)^{3}=\frac{(x_{n-2})^{4}}{(x_{n-3})^{3}}.$
-Iterating this process yields the following inequality:
+Iterating this process yields the following chain of inequalities:
 $x_{n+1}\leq\frac{(x_{n-1})^{3}}{(x_{n-2})^{2}}\leq\frac{(x_{n-2})^{4}}{(x_{n-3})^{3}}\leq\frac{(x_{n-3})^{5}}{(x_{n-4})^{4}}\leq\ldots\leq\frac{(x_{N+2})^{n-N}}{(x_{N+1})^{n-N-1}}\leq x_{N}\left(\frac{x_{N+1}}{x_{N}}\right)^{n+1-N}.$ (10)
@@
 $(x_{N+3})^{2}$ $\geq x_{N+2}x_{N+4}\geq\left((x_{N})^{\frac{1}{3}}(x_{N+3})^{\frac{2}{3}}\right)x_{N+4}\Rightarrow x_{N+3}\geq(x_{N})^{\frac{1}{4}}(x_{N+4})^{\frac{3}{4}}.$
-This process may continues recursively, and as $n$ gets large,
+This process may continue recursively, and as $n$ gets large,
 $x_{n}\geq\left(x_{N}\right)^{\frac{1}{n-N+1}}\left(x_{n+1}\right)^{\frac{n-N}{n-N+1}}.$ (12)
@@
 $=(n-N)\left(\frac{1}{n-N}-\frac{1}{n-N+m}\right)$
-$=\frac{m}{n-N+m}$
+$=\frac{m}{n-N+m}.$
 Thus, inequality (13) is reduced to:
@@
 for all $n>N$. This gives the left-hand side of (2).
-To prove the right hand side of Theorem 1.1, it is necessary to establish relations for all entries in the sequence, in particular, for $x_{0},x_{1},\dots,x_{N}$.
-To achieve this, define the sequence $\{a_{n}\}$ recursively by $a_{n}=x_{n}$ for all $n\geq N+1$ and $1\leq a_{n-1}\leq\frac{(a_{n})^{2}}{a_{n+1}}$ for $n=1,2,\ldots,N+1$.
-In particular, one may chose $a_{n}=x_{n}$ for all $n\geq N+1$ and for $n\leq N$,
-$a_{N}:=\frac{(a_{N+1})^{2}}{a_{N+2}}=\frac{(x_{N+1})^{2}}{x_{N+2}}$
-$a_{N-1}:=\frac{(a_{N})^{2}}{a_{N+1}}=\frac{\left(\frac{(x_{N+1})^{2}}{x_{N+2}}\right)^{2}}{x_{N+1}}=\frac{(x_{N+1})^{3}}{(x_{N+2})^{2}}$
+To prove the abundantness, it is necessary to establish relations for all entries in the sequence.
+To achieve this, construct a sequence $\{a_{n}\}$ recursively by $a_{n}=x_{n}$ for all $n\geq N+k$ and for $n<N+k$,
+$a_{N+k-1}:=\frac{(a_{N+k})^{2}}{a_{N+k+1}}=\frac{(x_{N+k})^{2}}{x_{N+k+1}}$ (15)
+$a_{N+k-2}:=\frac{(a_{N+k-1})^{2}}{a_{N+k}}=\frac{\left(\frac{(x_{N+k})^{2}}{x_{N+k+1}}\right)^{2}}{x_{N+k}}=\frac{(x_{N+k})^{3}}{(x_{N+k+1})^{2}}$
 $\vdots$
-$a_{N-N}=a_{0}:=\frac{(x_{N+1})^{N+2}}{(x_{N+2})^{N+1}}.$
-Hence, $(a_{n+1})^{2}\geq a_{n}a_{n+2}$ for all $n\geq 0$. A similar process to that outlined in (5) yields
+$a_{0}=\frac{(x_{N+k})^{N+k+1}}{(x_{N+k+1})^{N+k}}.$ (16)
+Hence, $(a_{n+1})^{2}\geq a_{n}a_{n+2}$ for all $n\geq 0$ and $a_{0}\geq 1$ by Condition (3). Now from Asai, Kubo, & Kuo [3, Theorem 2] or Gajdzica, Miska, & Ulas [18, Theorem 4.4], one obtains
+$a_{n}a_{m}\geq a_{n+m}\quad\text{for all}\quad m,n\geq 0.$ (17)
+Here, we present an alternative proof of (17) compared to the one given in [3, 18], as it aligns better with the approach used in the initial part of the proof.
+A similar process to that outlined in (10) yields
 $a_{n+1}\leq\frac{(a_{n-1})^{3}}{(a_{n-2})^{2}}\leq\frac{(a_{n-2})^{4}}{(a_{n-3})^{3}}\leq\frac{(a_{n-3})^{5}}{(a_{n-4})^{4}}\leq\ldots\leq\frac{(a_{2})^{n}}{(a_{1})^{n-1}}=a_{1}\left(\frac{a_{2}}{a_{1}}\right)^{n}.$ (18)
@@
 $(a_{n+m})^{\frac{1}{n+m-1}}>(a_{1})^{\frac{1}{n+m-1}+1}.$ (21)
-Note by Condition (2) that $a_{0}\geq 1$. And by log-concavity, $a_{1}^{2}\geq a_{0}a_{2}\geq a_{2}$ implying that $\frac{a_{2}}{a_{1}}\leq a_{1}$. Then from (10) it follows that
+Note by Conditions (3) and (16) that $a_{0}\geq 1$. And by log-concavity, $a_{1}^{2}\geq a_{0}a_{2}\geq a_{2}$ implying that $\frac{a_{2}}{a_{1}}\leq a_{1}$. Then from (18) it follows that
 $(a_{n+m})^{\frac{1}{n+m-1}}\leq\left[a_{1}\left(\frac{a_{2}}{a_{1}}\right)^{n+m-1}\right]^{\frac{1}{n+m-1}}$ $\leq\left(\frac{a_{2}}{a_{1}}\right)\left(a_{1}\right)^{\frac{1}{n+m-1}}$
 $\leq(a_{1})^{\frac{1}{n+m-1}+1}.$ (22)
-Hence
-Both (13) and (14) cannot be simultaneously true, this gives a contradiction. Hence $a_{m+n}\leq a_{m}a_{n}$ for all $m,n\geq 0$ and in particular $m,n\geq N+1$ completing the proof of Theorem 1.1. ∎
-Consider the consequences of starting this proof with the sequence $y_{n}=x_{n}^{-1}$.
-Then $y_{n}$ is log-convex for all $n>N$ and the same proof can be adapted where the only change is the reversal of every inequality sign, completing the proof of Theorem 1.2.
-∎
+Hence both (21) and (22) cannot be simultaneously true, this gives a contradiction. Hence $a_{m+n}\leq a_{m}a_{n}$ for all $m,n\geq 0$ and in particular $m,n\geq N+k$.
+Now, since $M\geq N+k\geq N$ then following the same steps of the proof of (14) we find
+$x_{n}\geq(x_{M})^{\frac{m}{n-M+m}}(x_{n+m})^{\frac{n-M}{n-M+m}}$
+and in particular
+$x_{M+1}\geq(x_{M})^{\frac{m}{m+1}}(x_{M+1+m})^{\frac{1}{m+1}}.$
+This gives
+$\frac{x_{M+1+m}}{x_{M+1}}\leq\left(\frac{x_{M+1}}{x_{M}}\right)^{m}.$
+Therefore by Condition (4)
+$\frac{x_{M+1+m}}{x_{M+1}}\leq x_{m}$
+for all $d\leq m\leq N+k$. Since the sequence $\{\frac{x_{n+1}}{x_{n}}\}_{n\geq N}$ is decreasing then
+$\frac{x_{n+m}}{x_{n}}\leq\frac{x_{M+1+m}}{x_{M+1}}\leq x_{m}$
+for all $n\geq M+1$ and $d\leq m\leq N+k$. This completes the proof of Theorem 1.1.
 3. Various Partition Functions
@@
 This section considers a few, well-studied sequences and partition functions and determines for which small indices abundance fails.
-3.1. The Distinct Partition Function, Partitions into Odd Parts
-The Distinct Partition Function $P_{d}(n)$ counts the number of ways to partition $n$ into distinct parts, where for all $i$ and $j$, $\lambda_{i}\neq\lambda_{j}$.
-$P_{d}(n)$ $=1,1,1,2,2,3,4,5,6,8,10,12,15,18,22,27,32,38,46,54,64,76,89,104,122,142,165,$
-[…]
-Curiously, Euler showed that $P_{d}(n)$ is equivalent to the number of partitions of $n$ using only odd integers [14, 25, 28].
-For example, $P_{d}(4)=2$ since there are only two ways to partition $4$ into distinct parts: $4=4=3+1$ (and only two ways to partition $4$ into odd parts: $4=3+1=1+1+1+1$).
-In 2023, Dong & Ji [11] proved that the sequence $\{P_{d}(n)\}$ is log-concave for all $n\geq 33$.
-Note that $P_{d}(n)$ satisfies Condition (2) for $n=33$: $448^{\frac{1}{33}}>512^{\frac{1}{34}}$.
-The result follows from Theorem 1.1 and numerical computation for small indices.
-Theorem 3.1.
-The distinct partition function satisfies the inequality $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$ where instances of equality and failure occur as given in Table 1.
-Table 1. Equality and Failure for $P_{d}(a)P_{d}(b)\geq P_{d}(a+b)$
-Equality at $(a,b)$ Failure at $(a,b)$
-$(1,1),(3,1),$ $a=1,2$
-$a=5,6,7,8\quad b=3,$ $(4,3),(5,5)$
-$a=15,16,17\quad b=4,$ $44\leq a\leq 14\quad b=$ 4
-$a=6,7,8\quad b=5$
-3.2. The Overpartition Function
-The overpartition function, $\overline{p}(n)$, counts the number of partitions of $n$ in which the last occurrence of a number can be distinguished. For example, $\overline{p}(4)=14$:
-$\ 1+1+1+1,\quad 1+1+1+\overline{1},\quad 1+1+2,\quad 1+1+\overline{2},\quad 1+\overline{1}+2,$
-$1+\overline{1}+\overline{2},\quad 2+2,\quad 2+\overline{2},\quad 1+3,\quad 1+\overline{3},\quad\overline{1}+3,\quad\overline{1}+\overline{3},\quad 4,\quad\overline{4}.$
-[…]
-In 2014, Engel [13] proved that $\{\overline{p}(n)\}$ function is log-concave for $n\geq 2$.
-Note that $\overline{p}$ satisfies Condition (2) for $n=2$: $\sqrt{4}\geq\sqrt[3]{8}$.
-The result follows from Theorem 1.1 and numerical computation for small indices.
-Theorem 3.2.
-The overpartition function satisfies the inequality $\overline{p}(a)\overline{p}(b)\geq\overline{p}(a+b)$ where instances of equality occur at $(1,1)$ and $(1,2)$.
-The following subsections, 3.3, 3.4, and 3.5, have been proven to satisfy log-concavity for sufficiently large $n$, however, the smallest such index $N$ such that each sequence is log-concave for all $n\geq N$ is either unknown or only conjectured.
-On the other hand, it is known that the asymptotics of these sequences are each of the form $n^{a}e^{c\cdot n^{b}}$ for real numbers $a,b,$ and $c$ where $0<b\leq 1$.
-This implies that the $n$th root of these sequences are each decreasing, hence, there is an index $N$ such that $x_{N+1}^{\frac{1}{N+1}}\geq x_{N+2}^{\frac{1}{N+2}}$.
-3.3. Fractional Partition Function
-Let $\alpha$ be a real number and define the fractional partition function $p_{\alpha}(n)$ by
-$\sum_{n=0}^{\infty}p_{\alpha}(n)q^{n}=(q;q)_{\infty}^{\alpha}$
-In 2019 Huat & Chan [6] studied the properties of $p_{\alpha}(n)$ where alpha is a rational number.
-The next year, Iskander, Jain & Talvola [23] proved that for any real $\alpha$ there exists an integer $N_{\alpha}$ such that the fractional partition function $p_{\alpha}(n)$ is log-concave $n>N_{\alpha}$.
-Theorem 3.3.
-There is an index $M_{\alpha}$ such that the sequence $p_{\alpha}(n)$ satisfies the inequality $p_{\alpha}(n)p_{\alpha}(m)>p_{\alpha}(n+m)$ for all $n,m>M_{\alpha}$
-3.4. The Multiset Partition Function
-Let $k$ be a fixed positive integer and let $\mathcal{A}=\{a_{i}\}_{i=1}^{\infty}$ be a non-decreasing sequence of positive integers and define a restricted partition of $n$ the sequence of positive integers $\lambda_{1},\lambda_{2},\ldots,\lambda_{j}$ such that $n=\lambda_{1}+\lambda_{2}+\ldots+\lambda_{j}$ where each $\lambda_{i}$ belongs to the multiset $\{a_{1},a_{2},...,a_{k}\}$.
-The multiset partition function $p_{\mathcal{A}}(n,k)$ counts all partitions of $n$ among the first $k$ elements of the multiset $\mathcal{A}$.
-The sequence $\{p_{\mathcal{A}}(n)\}$ was studied in 2022 by Gajdzica [15] who proved log-concavity under certain conditions of $n$ and $k$.
-This result was futher improved by Gajdzica, Heim, & Markus [16].
-Theorem 3.4.
-There exists a sufficiently large index such that the multiset partition function satisfies the inequality $p_{\mathcal{A}}(n)p_{\mathcal{A}}(m)\geq p_{\mathcal{A}}(n+m)$ under the same conditions Gajdzica gives in his Theorem 1.1 [15].
+The following subsections have been proven to satisfy log-concavity for sufficiently large $n$, however, the smallest such index $N$ such that each sequence is log-concave for all $n\geq N$ is either unknown or only conjectured.
+On the other hand, it is known that the asymptotics of these sequences are each of the form $n^{r}e^{t\cdot n^{s}}$ for real numbers $s,r,$ and $t$ where $0<s\leq 1$.
+This implies that for some $n\geq n_{0}$ the $n$th root of these sequences are each decreasing, hence, there is an index $N$ such that $x_{N+1}^{\frac{1}{N+1}}\geq x_{N+2}^{\frac{1}{N+2}}$.
 3.1. The $k$-th Power Partition Function
@@
 $1+1+1+1$
-Independently, both Benfield, Paul, & Roy [5] and Sullivan [32] proved that $p^{k}(n)$ is log-concave for all $k$ and sufficiently large $n$.
+Independently, both Benfield, Paul, & Roy [5] and O’Sullivan [35] proved that $p^{k}(n)$ is log-concave for all $k$ and sufficiently large $n$.
 Theorem 3.1.
-There exists an index $M_{k}$ such that for all $k$, the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ holds for $a,b>M_{k}+1$.
+There exists an index $N_{k}$ such that for all $k$, the inequality $p^{k}(m)p^{k}(n)\geq p^{k}(m+n)$ holds for $n\geq m\geq N_{k}$.
+Conjecture 3.2 gives instances $(m,n)$ where the inequality $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$ is equal and when it fails. The case $k=1$ was proven by Bessenrodt & Ono [33].
+Conjecture 3.2.
+The square power partition function satisfies the inequality $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$ with instances of failure or equality as given in Table $2$:
+Equality at $(m,n)$ with $n\geq m$ Failure at $(m,n)$ with $n\geq m$ $m=1$, $n=1,2,4,5,6,9,10,13,14,18,22$ $m=1$, $n\neq 1,2,4,5,6,9,10,13,14,18,22$ $m=2,n=4,5,9,13$ $m=2$, $n\neq 1,4,5,9,13$ $m=3$, $n=4$ $m=3$, $n\neq 1,2,4$ $m=4$, $n=5,6,7$ $m=5$, $n=5,6,8,11,15$ $m=5$, $n=7$ $m=6$, $n=8,10,12,14$ $m=6$, $n=6,7,11,15$ $m=7$, $n=8,9,12,13,19$ $m=7$, $n=7,10,11,14,15$
+Table 2. Equality and Failure for $p^{2}(m)p^{2}(n)\geq p^{2}(m+n)$
 What is the smallest index for which log-concavity fails? Define for $p^{k}(n)$, the smallest index $N_{k}$ such that for all $n>N_{k}$, the sequence ${p^{k}(n)}$ is log-concave.
@@
 The smallest $N_{k}$ such that $p^{k}(n)$ is log-concave for all $n>N_{k}$ for $k~{}\leq~{}6$ is given by:
-Table 2. Smallest $N_{k}$ for $k=1,\ldots,6$, OEIS: A346160 [37]
-$p^{k}(n)$ $p^{1}(n)$ $p^{2}(n)$ $p^{3}(n)$ $p^{4}(n)$ $p^{5}(n)$ $p^{6}(n)$
-$N_{k}$ $25$ $1041$ $15655$ $637854$ $2507860$ $35577568$
-Conjecture 3.7 gives instances $(a,b)$ where the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ is equal and when it fails for $k=1,2,3$.
-The case $k=1$ was proven by Bessenrodt & Ono [30].
-Note that equality always holds for $a+b<2^{k}$.
-Conjecture 3.7.
-The $k$-th power partition function satisfies the inequality $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$ with instances of failure or equality as given in Table 3:
-Table 3. Equality and Failure for $p^{k}(a)p^{k}(b)\geq p^{k}(a+b)$
-$p_{k}(n)$ Equality Failure
-$a<2^{1},\ a+b>2^{1}$
-$p_{1}(n)$ $(2,7),(2,6),(3,4)$ $a=2,3,4,5\quad b=2$
-$a=3,5\quad b=3$
-$a=5,6,7\quad b=4$ $a<2^{2},\ a+b>2^{2}$
-$p_{2}(n)$ $a=5,6,8,11,15\quad b=5$ $(7,5)$
-$a=8,10,12,14\quad b=6$ $a=6,7,11,15\quad b=6$
-$a=8,9,12,13,19\quad b=7$ $a=7,10,11,14,15\quad b=7$
-$a=15,23\quad b=9$ $a<2^{3},\ a+b>2^{3}$
-$p_{3}(n)$ $a=14,15,22,23\quad b=10$ $a=15,23\quad b=12$
-$a=13,14,15,21,22,23\quad b=11$ $a=14,15,22,23,b=13$
-$a=12,13,14,20,21,22\quad b=12$ $a=14,15,21,22,23\quad b=14$
-$a=13,19,20,21,23\quad b=13$ $a=15,20,21,22,23\quad b=15$
-$a=18,19,20,26\quad b=14$
-$a=17,18,19,26\quad b=15$
-$(23,20),(22,21),(23,21),(22,22),(23,22),(23,23)$
+Table 3. Smallest $N_{k}$ for $k=2,\ldots,6$, OEIS: A346160 [40]
+$p^{k}(n)$ $p^{2}(n)$ $p^{3}(n)$ $p^{4}(n)$ $p^{5}(n)$ $p^{6}(n)$
+$N_{k}$ $1041$ $15655$ $637854$ $2507860$ $35577568$
+3.2. Fractional Partition Function
+Let $\alpha$ be a real number and define the fractional partition function $p_{\alpha}(n)$ by
+$\sum_{n=0}^{\infty}p_{\alpha}(n)q^{n}=(q;q)_{\infty}^{\alpha}$
+In 2019 Huat & Chan [8] studied the properties of $p_{\alpha}(n)$ where $\alpha$ is a rational number.
+The next year, Iskander, Jain, & Talvola [25] proved that for any real $\alpha$ there exists an integer $N_{\alpha}$ such that the fractional partition function $p_{\alpha}(n)$ is log-concave $n>N_{\alpha}$.
+Theorem 3.4.
+Let $\alpha>0$. Then there is an index $M_{\alpha}\geq N_{\alpha}$ such that the sequence $p_{\alpha}(n)$ satisfies the inequality $p_{\alpha}(n)p_{\alpha}(m)>p_{\alpha}(n+m)$ for all $n,m>M_{\alpha}$
+3.3. The Multiset Partition Function
+Let $k$ be a fixed positive integer and let $\mathcal{A}=\{a_{i}\}_{i=1}^{\infty}$ be a non-decreasing sequence of positive integers and define a restricted partition of $n$ as the sequence of positive integers $\lambda_{1},\lambda_{2},\ldots,\lambda_{j}$ such that $n=\lambda_{1}+\lambda_{2}+\ldots+\lambda_{j}$ where each $\lambda_{i}$ belongs to the multiset $\{a_{1},a_{2},...,a_{k}\}$.
+The multiset partition function $p_{\mathcal{A}}(n,k)$ counts all partitions of $n$ among the first $k$ elements of the multiset $\mathcal{A}$.
+The sequence $\{p_{\mathcal{A}}(n,k)\}$ was studied in 2022 by Gajdzica [16] who proved log-concavity under certain conditions of $n$ and $k$.
+This result was further improved by Gajdzica, Heim, & Neuhauser [17].
+Theorem 3.5.
+There exists a sufficiently large index such that the multiset partition function satisfies the inequality $p_{\mathcal{A}}(n,k)p_{\mathcal{A}}(m,k)\geq p_{\mathcal{A}}(n+m,k)$ under the same conditions Gajdzica gives in his Theorem 1.1 [16].
 4. Log-polynomial Sequences
-In a more general setting, Sullivan [33] made interesting connections between the data of log-polynomial sequences and log-concavity/log-convexity.
-A sequence $\alpha(n)$ is log-polynomial of degree $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ for $\kappa=\pm 1$ if there exists $\delta(n)>0$, $A(n)$, and $g_{k}(n)$ for $k=3,4,\ldots,m$ such that
-$\log{\left(\frac{\alpha(n+j)}{\alpha(n)}\right)}=A(n)j+\kappa\cdot\delta(n)^{2}j^{2}+\sum_{k=3}^{m}g_{k}(n)j^{k}+o\left(\delta(n)^{m+1}\right)$
+In a more general setting, O’Sullivan [36] made interesting connections between the data of log-polynomial sequences and log-concavity/log-convexity.
+A sequence $a(n)$ is log-polynomial of degree $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ for $\kappa=\pm 1$ if there exists $\delta(n)>0$, $A(n)$, and $g_{k}(n)$ for $k=3,4,\ldots,m$ such that
+$\log{\left(\frac{a(n+j)}{a(n)}\right)}=A(n)j+\kappa\cdot\delta(n)^{2}j^{2}+\sum_{k=3}^{m}g_{k}(n)j^{k}+o\left(\delta(n)^{m+1}\right)$
 for $j=1,2,\ldots,m+1$ where both $\delta(n)\rightarrow 0$ and $g_{k}(n)=o\left(\delta(n)^{k}\right)$ as $n\rightarrow\infty$.
-Sullivan [33] showed that if $\alpha(n)$ is log-polynomial for degrees $m\geq 2$, then $\alpha(n)$ is log-concave for large $n$ whenever $\kappa=-1$ and is log-convex for large $n$ whenever $\kappa=1$.
-In particular, Sullivan showed [33] that if $\alpha(n)=n^{a}e^{c\cdot n^{b}}$ for real numbers $a$, $b$, and $c$ where $a$ and $c$ are not both zero, $20<b<$ 2, and $b(b-1)c\neq 0$, then $\alpha(n)$ is log-polynomial with data $\{A(n),\kappa,\delta(n)\}$ where
-$A(n)=-\frac{a}{n}+bc\cdot n^{b-1}\quad\ \quad\kappa=\text{sgn}(b(b-1)c)\quad\ \quad\kappa\cdot\delta(n)^{2}=-\frac{a}{2n^{2}}+\frac{b(b-1)c}{2n^{2-b}}$
+O’Sullivan [36] showed that if $a(n)$ is log-polynomial for degrees $m\geq 2$, then $a(n)$ is log-concave for large $n$ whenever $\kappa=-1$ and is log-convex for large $n$ whenever $\kappa=1$.
+In particular, he showed [36] that if $a(n)=n^{r}e^{t\cdot n^{s}}$ for real numbers $r$, $s$, and $t$ where $r$ and $t$ are not both zero, $20<s<$ 2, and $s(s-1)t\neq 0$, then $a(n)$ is log-polynomial with data $\{A(n),\kappa,\delta(n)\}$ where
+$A(n)=-\frac{r}{n}+st\cdot n^{s-1}\quad\ \quad\kappa=\text{sgn}(s(s-1)t)\quad\ \quad\kappa\cdot\delta(n)^{2}=-\frac{r}{2n^{2}}+\frac{s(s-1)t}{2n^{2-s}}$
 for all degrees $m\geq 2$.
 Theorem 4.1.
-If a sequence $\alpha(n)$ is log-polynomial for degrees $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ and $0<b\leq 1$, then for sufficiently large integers $n$ and $m$, the sequence satisfies $\alpha(n)\alpha(m)\geq\alpha(n+m)$ whenever $\kappa=-1$ and satisfies $\alpha(n)\alpha(m)\leq\alpha(n+m)$ whenever $\kappa=1$ with finitely many instances of equality or failure.
+If a sequence $a(n)$ is log-polynomial for degrees $m\geq 2$ with data $\{A(n),\kappa,\delta(n)\}$ and $0<s\leq 1$, then for sufficiently large integers $n$ and $m$, the sequence satisfies $a(n)a(m)\geq a(n+m)$ whenever $\kappa=-1$ and satisfies $a(n)a(m)\leq a(n+m)$ whenever $\kappa=1$ with finitely many instances of equality or failure.
 5. Sufficient But Not Necessary Condition for Partition Functions
@@
 $2^{0}+2^{0}+2^{0}+2^{0}$
-In 2019, Ganguly [18] showed that $b^{m}(n)$ is neither log-concave nor log-convex.
-Yet, for each $m$, the sequence $\{b^{m}(n)\}_{n\in\mathbb{N}}$ appears to be abundant with only finitely many instances of equality or failure.
+In 2019, Ganguly [19] showed that $b^{m}(\alpha)$ is neither log-concave nor log-convex.
+Yet, for each $m$, the sequence $\{b^{m}(\alpha)\}_{\alpha\in\mathbb{N}}$ appears to be abundant with instances of equality or failure according to congruence classes for $\alpha<m$ and only finitely many instances of equality (and no instances of failure) for $\alpha\geq m$.
+A regular pattern appears for $m\geq 6$.
 Conjecture 5.1.
-For each $m$, the sequence given by $\{b^{m}(n)\}_{n\in\mathbb{N}}$ satisfies the inequality $b^{m}(\alpha)b^{m}(\beta)~{}\geq~{}b^{m}(\alpha+\beta)$ with instances of equality or failure as given by Table 4.
+For each $2\leq m\leq 5$, the sequence given by $\{b^{m}(\alpha)\}_{\alpha\in\mathbb{N}}$ satisfies the inequality $b^{m}(\alpha)b^{m}(\beta)~{}\geq~{}b^{m}(\alpha+\beta)$ with instances of equality or failure as given by Table 4.
+[…]
 Table 4. Equality and Failure for $b^{m}(\alpha)b^{m}(\beta)\geq b^{m}(\alpha+\beta)$
-$b^{m}(n)$ Equality at $(a,b)$ Failure at $(\alpha,\beta)$
-$b^{2}(n)$ $(2,2),(3,2),(3,9)$ $a,b\leq 1$,
-$(3,3),(5,3),(7,3)$
-$b^{3}(n)$ $(7,8),(8,8)$ $a,b\leq 2$,
-$(4,5),(4,8),(5,5),(5,7),(5,8)$
-$b^{4}(n)$ $(5,7),(5,11),(5,15),(6,6)$ $a,b\leq 3$,
-$(6,7),(6,10),(6,11),(6,14)$
-$(6,15),(7,7),(7,9),(7,10)$
-$(7,11),(7,13),(7,14),(7,15)$
-And for $m\geq 5$, Failure only occurs for $\alpha,\beta\leq m$.
-For $m=5$, equality occurs for pairs $(\alpha,\beta)$ where $12m-2\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$.
-For $m\geq 6$, equality occurs for pairs $(\alpha,\beta)$ where $12m-3\leq\alpha\leq 2m-$ 1 and $m+1\leq\beta\leq 2m-1$.
-For example,
-$b^{5}(n)\ \text{has equality at}$ $b^{6}(n)\ \text{has equality at}$
-$(8,7),(8,8)$ $(9,9)$
-$(9,6),(9,7),(9,8),(9,9)$ $(10,8),(10,9),(10,10)$
-$(11,7),(11,8),(11,9),(11,10),(11,11)$
 6. Final Thoughts
@@
 Acknowledgements
-The authors would like to extend sincere gratitude to the anonymous referee, whose keen observations and criticism throughout the review process has greatly improved the quality of this paper.
+The author would like to extend sincere gratitude to the anonymous referee, whose keen observations and criticism throughout the review process has greatly improved the quality of this paper.